OpenStudy (anonymous):

the curve y=ax^3 +bx passes throught the point (1,7). thr tangent at this point is parallel to the line y=2x-6

OpenStudy (dumbcow):

slope of tangent line is dy/dx since it is parallel to y=2x-6, the slope must be 2 --> dy/dx = 2 --> 3ax^2+b = 2 x=1 --> 3a +b = 2 not sure what they want from here..

OpenStudy (anonymous):

Not sure you even had to take the derivative because they give not only a slope but a point the line passes through.

OpenStudy (dumbcow):

oh right :{ so equation of tangent line is y-7 = 2(x-1)

OpenStudy (anonymous):

Yeah, I was waiting to see if sarah returned before answering, but she flew the coop as I was starting to reply and so I just stopped....lol

OpenStudy (dumbcow):

oh but if you need to solve for a,b you can use my prev post plus fact that if you plug in point you get 7 = a+b

OpenStudy (anonymous):

I considered that but there was no directions when I got here, so I was going ask Sara what was the question wanting.?

OpenStudy (dumbcow):

hmm yes the entire question wasn't posted ... i usually speculate there are only so many unknowns to find :)

OpenStudy (anonymous):

Ah well, guess she can post a reply if she returns.

OpenStudy (anonymous):

I had a question the other day about finding the value of k in a problem that had no k to be found....after a series of questions, I was able to determine that they were comparing the equation to a standard form of a square root, and k was part of the vertex...lol.

OpenStudy (anonymous):

hehe sorry they say find a and b

OpenStudy (anonymous):

From what @dumbcow has shown, we get the system of equations: 3a + b = 2 a + b = 7 Now one can simply solve for a and b. I will use elimination to solve for a and b. 3a + b = 2 - a + b = 7 ------------ 2a = -5 a = -5/2 Now plug in this value of a back in to one of the equations and solve for b. (-5/2) + b = 7 b = 7 + 5/2 b = 14/2 + 5/2 b = 19/2 Therefore, a = -5/2 and b = 19/2 @sarahlovesyumn

OpenStudy (anonymous):

thank u :) but how did u get 3a+b=2?

OpenStudy (dumbcow):

@sarahlovesyumn, refer to very first post at top of page, i explain it there :)

OpenStudy (dumbcow):

to clarify more.... the derivative of function when x=1 is "3a+b"